by Charpits AE sobylve z=p^2x+q^2y
1
Expert's answer
2019-10-22T07:58:56-0400

We rewrite this equation in the form


"z-p^2x-q^2y=0"



which allows us to highlight the function "F(x,y,z,p,q)=z-p^2x-q^2y" , which we need for Charpit's Method.


Charpit's equation is:



"\frac{dx}{F_p}=\frac{dy}{F_q}=\frac{dz}{pF_p+qF_q}=\frac{dp}{-\left(F_x+pF_z\right)}=
\frac{dq}{-\left(F_y+qF_z\right)}"

In our case,


"F(x,y,z,p,q)=z-p^2x-q^2y\longrightarrow\\[0.3cm]
F_p=-2px\quad F_q=-2qy\\[0.3cm]
pF_p+qF_q=-2p^2x-2q^2x=-2\left(p^2x+q^2y\right)\equiv-2z\\[0.3cm]
F_x+pF_z=-p^2+p\quad F_y+qF_z=-q^2+q"

Then, our equation takes the form



"\frac{dx}{-2px}=\frac{dy}{-2qy}=\frac{dz}{-2z}=\frac{dp}{p^2-p}=
\frac{dq}{q^2-q}"



1) We are interested in the first and fourth element:


"\frac{dx}{-2px}=\frac{dp}{p^2-p}\longrightarrow\frac{dx}{-2px}=\frac{dp}{p(p-1)}\longrightarrow\frac{dx}{-2x}=\frac{dp}{(p-1)}\\[0.3cm]
-\frac{1}{2}\cdot\int\frac{dx}{x}=\int\frac{dp}{p-1}\longrightarrow\\[0.3cm]
-\frac{1}{2}\ln|x|=\ln|p-1|-\ln|A|\longrightarrow
\ln\left|\frac{A}{\sqrt{x}}\right|=\ln|(p-1)|\longrightarrow\\[0.3cm]
\frac{A}{\sqrt{x}}=p-1\longrightarrow\boxed{p=\frac{A+\sqrt{x}}{\sqrt{x}}}"

2) We are interested in the second and fifth element:



"\frac{dy}{-2qy}=\frac{dq}{q^2-q}\longrightarrow\frac{dy}{-2qy}=\frac{dq}{q(q-1)}\longrightarrow\frac{dy}{-2y}=\frac{dq}{(q-1)}\\[0.3cm]
-\frac{1}{2}\cdot\int\frac{dy}{y}=\int\frac{dq}{q-1}\longrightarrow\\[0.3cm]
-\frac{1}{2}\ln|y|=\ln|q-1|-\ln|B|\longrightarrow
\ln\left|\frac{B}{\sqrt{y}}\right|=\ln|(q-1)|\longrightarrow\\[0.3cm]
\frac{B}{\sqrt{y}}=q-1\longrightarrow\boxed{q=\frac{B+\sqrt{y}}{\sqrt{y}}}"

It remains to substitute the obtained results in the initial equation:



"z=p^2x+q^2y\longrightarrow z=\left(\frac{A+\sqrt{x}}{\sqrt{x}}\right)^2x+\left(\frac{B+\sqrt{y}}{\sqrt{y}}\right)^2y\\[0.3cm]
z=\frac{A^2+2A\sqrt{x}+x}{x}\cdot x+\frac{B^2+2B\sqrt{y}+y}{y}\cdot y\\[0.3cm]
\boxed{z=x+y+2\left(A\sqrt{x}+B\sqrt{y}\right)+A^2+B^2}"

ANSWER



"z=x+y+2\left(A\sqrt{x}+B\sqrt{y}\right)+A^2+B^2"